OpenStudy (anonymous):

-3x^3+x^2+2x rises as x grows very large true or false

OpenStudy (mathmale):

We often discuss "dominant term" in situations such as this one. In your polynomial, f(x)=-3x^3+x^2+2x, the dominant term is that involving the largest power of x, i. e., x^3. Ask yourself: How does y=x^3 behave as x continues to increase? Ask you rself: How does y=-x^3 behave?

OpenStudy (anonymous):

im not sure....

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Have you tried graphing these two functions? They are extremely common ones, so you should pretty much know their shapes by heart. if you don't, couldn't you use a graphing calculator to determine the shapes of both?

OpenStudy (anonymous):

I did graph it and I think its true?